### v3

@@ 4.6 Computational Setup and Running Times
All experiments were conducted on the Charité High-Performance Cluster using Nvidia A100 GPUs with 80 GB memory, configured with one, four, or eight GPUs (DGX systems). Running the GTE-Qwen2-7B model with our serialization truncated at 4,096 tokens on an 8-GPU DGX system required approximately 20 hours. For the LLM2Vec-Llama-3.1-8B model, runtime errors occurred during multi-GPU experiments with the full dataset. These issues were resolved by splitting the data into smaller batches, which introduced additional overhead. Additionally, we optimized the LLM2Vec code by removing an initial text pruning step that took approximately eight hours when using the full data. Using this modified setup, the LLM2Vec-Llama-3.1-8B model took approximately 30 hours to complete on eight A100 GPUs. The calculation of the embeddings for the 387,464 UKB patients took approximately 35 hours on a single GPU per task for the LLM-based methods. For CLMBR-T-Base, the calculation was faster, only requiring around eight hours.
@@ 4.8 External Validation on UK Biobank
External validation was performed using data from the UKB, a large-scale prospective cohort study comprising 502,489 UK participants recruited between 2006 and 2010, with a median follow-up of 13.8 years. We used linked EHR data from primary care (General Practitioner, GP) and secondary care (Hospital Episode Statistics, HES), providing information on diagnoses, procedures, and prescriptions.
@@
Initial data preprocessing, including cleaning, feature extraction, missing value imputation, and endpoint selection followed the methodology described in [34]. All health records were mapped to the OMOP CDM using mapping tables provided by the UKB, SNOMED International, and the OHDSI community for mapping concepts from the provider and country-specific non-standard vocabularies to OMOP standard vocabularies. Participants lacking any recorded GP or HES events either before or after their recruitment date were excluded, resulting in a validation cohort of 387,464 individuals. Diagnostic codes were mapped to Phecodes X [53, 54] (derived directly from source ICD-10 or by mapping from SNOMED to ICD-10 codes and subsequently to Phecodes X) primarily for standardized endpoint definition and cohort selection. To avoid redundancy with source codes used as features, the Phecodes themselves were excluded during the creation of patient sequences for model input. Due to significant challenges in mapping and harmonizing UKB laboratory values [55], and to ensure comparability across models and tasks, laboratory data were excluded entirely, differing from the use of binary labels in [34]. The final feature set comprised conditions (SNOMED, CVX), medications (RxNorm), and procedures (SNOMED).
@@
Given that UKB represents a general population cohort, we adjusted the proposed prediction tasks to define relevant longitudinal health trajectory tasks. These tasks included: (1) prediction of all-cause hospitalization within the next year (operational outcomes), (2) prediction of incident diagnoses for a set of selected conditions (assignment of new diagnoses), and (3) prediction of all-cause mortality (mortality prediction). The selection of incident diseases for the assignment of new diagnoses task group largely followed [34], focusing on common conditions, diseases lacking established risk stratification tools, and specific cardiovascular conditions. From the initial 24 endpoints proposed in [34], we treated all-cause death as a separate task. For the assignment of new diagnoses and mortality tasks, patients with a diagnosis of the respective endpoint recorded prior to their UKB recruitment date were excluded to prevent data leakage; this exclusion was not applied to the hospitalization task due to high incidence. Following task-specific exclusions, the final sample size available for analysis ranged from 300,344 to 359,250 patients, depending on the prediction task (see Table 6 for details). For all three task groups, the prediction window was set to one year after the prediction year.
@@
The pre-trained CLMBR-T-Base model operates with a fixed vocabulary (65,536 total, 26,249 unique codes). To use this model, we mapped the 50,702 unique medical codes (SNOMED CT, RxNorm, CVX) present in our processed UKB cohort to the CLMBR-T-Base vocabulary. This mapping followed the steps similar to those implemented in the FEMR package101010https://github.com/som-shahlab/femr/tree/0ebf454303091fa83c9e4023e7db9ba9b68af09a, involving direct code matching where possible, supplemented by indirect mapping via the OHDSI ATHENA vocabulary using ’Maps to’ relationships and inclusion of ancestor concepts. This process successfully mapped 7,969 unique UKB codes to the CLMBR-T-Base vocabulary (in format ontology/code), which means that the remaining codes were implicitly excluded from the input sequences for this model. After creating patient timelines, adding information about birth date, ethnicity, and visits, the data was converted into the MEDS standard [56], from which the embeddings were calculated. Additionally, UKB ethnicities were converted to the ethnicity groups used by CLMBR-T-Base. For transforming the patient information into embeddings, we mainly followed the code provided by FEMR (mainly the convert_patient function111111https://huggingface.co/StanfordShahLab/clmbr-t-base) with minor modifications to enable batch processing.
@@
Similar to the approach of Wornow et al. [31], we performed an ontology expansion to improve the counts baseline by adding direct parent and grandparent concepts. This increased the unique code count from 51,677 to 69,850, aiming to capture broader hierarchical relationships between medical concepts. However, the resulting high dimensionality posed memory challenges for a standard counts matrix. To mitigate this, we implemented feature selection by removing codes present in less than 50 individuals (0.01% of the cohort), reducing the feature space to 24,509 unique codes. To ensure a more comparable baseline to the other models, which implicitly incorporate patient context, we explicitly added normalized age at prediction time and coded sex as additional features to the feature matrix used for the count-based baseline.
@@
Similar to before, all approaches were evaluated using a five-fold cross-validation. Within each fold, hyperparameters were optimized on a dedicated validation set of similar sample size to that of the training partition in that fold. For each fold, patients were randomly assigned to train, validation, and test sets. The test set for each task within a fold was constructed by selecting all available positive cases (capped to 10,000), and an equal number of negative cases after allocating patients to the train/validation sets. To specifically assess performance in low-data scenarios, mimicking potential clinical applications with limited examples, we again evaluated models using incrementally larger subsets of the training data, with the final evaluation point capped at a sample size of 256 patients from the train set, even when larger test sets were available.
4.9 Ablations of EHR Serialization
To better understand the contribution of various components in the EHR serialization process to the performance of the LLM embedding models, we conducted a series of ablation experiments. First, we assessed the impact of task-specific instructions by replacing them with a generic instruction (Table 8) or removing them entirely to determine their influence on the final embeddings. Subsequently, we performed additional ablations by systematically excluding specific components of the serialization. These included demographic data, aggregated LOINC codes, and both visit summaries and detailed entries. Furthermore, we examined the effect of removing specific fields from the detailed visit entries, such as conditions, medications, and procedures. Throughout these experiments, the rest of the pipeline was kept consistent to isolate the effects of the removed components. This approach allowed us to identify which parts of the serialization process were most critical for generating effective EHR representations, providing insights into how these models use structured medical data for prediction tasks.
@@ 4.10 Effect of Different Time Windows
To examine the influence of recency on predictive performance, we varied the time window preceding the prediction date used during EHR text serialization. We evaluated GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B across six intervals: one day, one week (7 days), one month (30 days), one year (365 days), three years (1,095 days), and full history. For each window, only events occurring within the specified interval before the prediction time were included. Hence, only data from the respective time window contributed to the aggregated information and the visit data in the EHR serialization. All other aspects of the serialization, including structure, formatting, and instruction prompts remained unchanged to isolate the effect of the temporal window.
@@ 4.11 Effect of Different Context Sizes
We investigated the impact of varying context sizes in the LLM embedding models. We wanted to determine whether encoding information from older visits enhances prediction performance and whether longer inputs might dilute critical details, such as laboratory values, in the final embeddings. Specifically, we evaluated GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B models with input token limits of 512, 1,024, 2,048, and 8,192 tokens. Input tokens exceeding these thresholds were discarded. Due to the design of our EHR serialization process, additional input tokens primarily consisted of medical concepts from past visits. By testing these varying context sizes, we aimed to assess the balance between capturing historical medical data and preserving the clarity of high-priority information within the embeddings.
4.12 Effect of Chunked Contexts
To further explore whether LLM embedding models effectively interpret their entire input, we compared the performance of models processing complete inputs versus segmented (chunked) inputs. For this, the 4,096-token inputs were divided into smaller chunks of sizes 512, 1,024, and 2,048 tokens121212In our implementation, we used these chunk sizes decreased by eight tokens to cater for special tokens that might be added by the implementations of the language models.. Separate embeddings were generated for each chunk, and a final embedding was obtained by averaging the embeddings of all chunks. This approach aligns with the mean pooling applied to the last layer that we used to compute embeddings. In this setup, task-specific instructions, placed at the beginning of the inputs, were included only in the first chunk. This design allowed us to evaluate the models’ ability to contextualize task instructions and whether chunking affects the overall performance.
@@ A.1 Additional Experimental Details
A.2 Effect of Different Time Windows
To assess how the recency of medical history affects prediction performance, we varied the temporal window preceding the prediction time during EHR text serialization. The results are summarized in Fig. 6 (left) and detailed in Table 9. GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B exhibited similar overall trends, though GTE-Qwen2-7B performed better with larger time windows. The one-week and one-month windows consistently yielded the highest AUROC scores, suggesting that recent clinical data carries the most predictive value for downstream tasks. Notably, the 24-hour window produced better results than using the full historical record. This trend suggests that recent clinical data often holds the most predictive value, while older information may introduce irrelevant or outdated signals. Among the two models, GTE-Qwen2-7B demonstrated stronger performance with longer time windows, indicating a higher capacity to extract relevant features from extended clinical narratives.
A.3 Effect of Different Context Sizes
We further tested the impact of varying context sizes on the performance of the LLM embedding models. As summarized in Fig. 6 (right) and Table 10, the two models exhibited distinct behaviors. For GTE-Qwen2-7B, the best overall performance was achieved with a context of 4,096 tokens, with only a slight decline observed when using 2,048 or 1,024 tokens. This suggests that the most relevant information is contained within the first 1,024 tokens and that the model can accurately extract these details even when presented within a longer context. In contrast, LLM2Vec-Llama-3.1-8B showed a marked improvement with a context size of 2,048 tokens, indicating that the model struggles to effectively handle longer sequences up to 4,096 tokens. Both models experienced a significant drop in performance when limited to only 512 tokens, highlighting that such a short context omits crucial information; a noteworthy point given that many existing language models are constrained to 512 input tokens [37]. Furthermore, extending the context to 8,192 tokens resulted in decreased performance for both models, implying limitations in extracting relevant information from very long inputs. This decline may be partially attributed to the mean pooling of the last hidden layers used for generating embeddings, which can dilute the impact of important token-level information; an alternative strategy, such as incorporating an additional attention mechanism, might help mitigate this issue.
A.4 Effect of Chunked Contexts
To investigate whether the models can process the full 4,096-token context cohesively, we conducted an experiment in which the serialized EHR input was divided into chunks of 512, 1,024, and 2,048 tokens. For each chunk, separate embeddings were generated and then averaged to create a final representation (see Table 11). For GTE-Qwen2-7B, the performance decrease with smaller chunks was relatively modest, indicating that the information contained within the full 4,096-token input remains effectively used even when segmented. Notably, using 512-token chunks yielded an overall AUROC performance of 0.731 (0.706-0.756), compared to 0.671 (0.647-0.695) when processing a contiguous 512-token context, which suggests that chunking can mitigate input constraints. In contrast, LLM2Vec-Llama-3.1-8B demonstrated improved performance with chunked inputs, consistent with its behavior on shorter context sizes. This enhancement was primarily driven by better lab value prediction, implying that LLM2Vec-Llama-3.1-8B is particularly effective when processing inputs of up to 2,048 tokens.
A.5 Detailed Ablation Results for EHRSHOT
@@
| GTE-Qwen2-7B | |||||
@@
| GTE-Qwen2-7B | |||||
| 8.192 context size | |||||
| 4.096 context size | |||||
| 2.048 context size | |||||
| 1.024 context size | |||||
| 512 context size | |||||
| LLM2Vec-Llama-3.1-8B | |||||
| 8.192 context size | |||||
| 4.096 context size | |||||
| 2.048 context size | |||||
| 1.024 context size | |||||
| 512 context size | |||||
| GTE-Qwen2-7B | |||||
| 1 x 4.096 token chunks | |||||
| 2 x 2.048 tokens chunks | |||||
| 4 x 1.024 tokens chunks | |||||
| 8 x 512 tokens chunks | |||||
| LLM2Vec-Llama-3.1-8B | |||||
| 1 x 4.096 token chunks | |||||
| 2 x 2.048 tokens chunks | |||||
| 4 x 1.024 tokens chunks | |||||
| 8 x 512 tokens chunks | |||||
A.6 Additional Performance Results on EHRSHOT
A.7 Additional Performance Results on UK Biobank

### v4

@@ 4.8 Computational Setup and Running Times
All experiments were conducted on the Charité High-Performance Cluster using Nvidia A100 GPUs with 80 GB memory, configured with one, four, or eight GPUs (DGX systems). Running the Qwen3-Emb-8B model with our serialization truncated at 8,192 tokens on an 8-GPU DGX system required approximately 40 hours. For the same setup Qwen2-Emb-7B required approximately 31 hours. For the Llama3.1-LLM2Vec-8B model, runtime errors occurred during multi-GPU experiments with the full dataset. These issues were resolved by splitting the data into smaller batches, which introduced additional overhead. Additionally, we optimized the LLM2Vec code by removing an initial word-boundary token-limit routine that iteratively re-tokenizes to ensure the final cut occurs at a word boundary; this step took approximately eight hours when using the full data. With our workaround, inputs are simply truncated at the maximum token number (i.e., the last token may correspond to an incomplete word), which matches how the Qwen implementation handles over-length inputs. Importantly, this modification does not perform any semantic text pruning or content selection and should have only a minimal effect on performance, limited to a potentially incomplete trailing word. Using this modified setup, the Llama3.1-LLM2Vec-8B model took approximately 46 hours to complete on eight A100 GPUs. The calculation of the embeddings for the 387,464 UKB patients took approximately 80 hours on a single GPU per task for Qwen2-Emb-7B and Llama3.1-LLM2Vec-8B and approximately 100 hours for Qwen3-Emb-8B, all for a token length of 8,192. For significantly reducing the overall runtime, the calculations can be split onto separate GPUs. Additionally, since the embedding for the disease prediction task is not disease specific, for each model only three embeddings had to be calculated per patient.
@@ 4.10 External Validation on UK Biobank
External validation was performed using data from the UKB, a large-scale prospective cohort study comprising 502,489 UK participants recruited between 2006 and 2010, with a median follow-up of 13.8 years. We used linked EHR data from primary care (General Practitioner (GP)) and secondary care (Hospital Episode Statistics (HES)), providing information on diagnoses, procedures, and prescriptions.
@@
Initial data preprocessing, including cleaning, feature extraction, missing-value imputation, and endpoint selection, followed the methodology described in [45]. All health records were mapped to the OMOP CDM using mapping tables provided by the UKB, SNOMED International, and the OHDSI community for mapping concepts from provider and country-specific non-standard vocabularies to OMOP standard vocabularies. Participants lacking any recorded GP or HES events either before or after their recruitment date were excluded, resulting in a validation cohort of 387,464 individuals. Diagnostic codes were mapped to Phecodes X [66, 67] (derived directly from source ICD-10 or by mapping from SNOMED to ICD-10 codes and subsequently to Phecodes X) primarily for standardized endpoint definition and cohort selection. To avoid redundancy with source codes used as features, the Phecodes themselves were excluded during the creation of patient sequences for model input. Due to significant challenges in mapping and harmonizing UKB laboratory values [68], and to ensure comparability across models and tasks, laboratory data were excluded entirely, differing from the use of binary labels in [45]. The final feature set comprised conditions (SNOMED, CVX), medications (RxNorm), and procedures (SNOMED).
@@
Given that UKB represents a general population cohort, we adjusted the proposed prediction tasks to define relevant longitudinal health trajectory tasks. These tasks included: (1) prediction of all-cause hospitalization within the next year (operational outcomes), (2) prediction of incident diagnoses for a set of selected conditions (assignment of new diagnoses), and (3) prediction of all-cause mortality (mortality prediction). The selection of incident diseases for the assignment of new diagnoses task group largely followed [45], focusing on common conditions, diseases lacking established risk stratification tools, and specific cardiovascular conditions. From the initial 24 endpoints proposed in [45], we treated all-cause mortality as a separate task. For the assignment of new diagnoses and mortality tasks, patients with a diagnosis of the respective endpoint recorded prior to their UKB recruitment date were excluded to prevent data leakage; this exclusion was not applied to the hospitalization task due to high incidence. Following task-specific exclusions, the final sample size available for analysis ranged from 300,344 to 359,250 patients, depending on the prediction task (see Table 7 for details). For all three task groups, the prediction window was set to one year after the prediction year.
@@
The pretrained CLMBR-T-Base model operates with a fixed vocabulary of 26,249 unique codes. To use this model, we mapped the 50,702 unique medical codes (SNOMED CT, RxNorm, CVX) present in our processed UKB cohort to the CLMBR-T-Base vocabulary. This mapping followed steps similar to those implemented in the FEMR package111111https://github.com/som-shahlab/femr/tree/0ebf454303091fa83c9e4023e7db9ba9b68af09a, involving direct code matching where possible, supplemented by indirect mapping via the OHDSI ATHENA vocabulary using ’Maps to’ relationships and inclusion of ancestor concepts. This process successfully mapped 7,969 unique UKB codes to the CLMBR-T-Base vocabulary (in format ontology/code), which means that the remaining codes were implicitly excluded from the input sequences for this model. After creating patient timelines, adding information about birth date, ethnicity, and visits, the data was converted into the MEDS standard [69], from which the embeddings were calculated. Additionally, UKB ethnicities were converted to the ethnicity groups used by CLMBR-T-Base. For transforming the patient information into embeddings, we mainly followed the code provided by FEMR (primarily the convert_patient function121212https://huggingface.co/StanfordShahLab/clmbr-t-base) with minor modifications to enable batch processing.
@@
Similar to the approach of Wornow et al. [44], we performed an ontology expansion to improve the count-based baseline by adding direct parent and grandparent concepts. This increased the unique code count from 51,677 to 69,850, aiming to capture broader hierarchical relationships between medical concepts. However, the resulting high dimensionality posed memory challenges for a standard counts matrix. To mitigate this, we implemented feature selection by removing codes present in fewer than 50 individuals (<0.01% of the cohort), reducing the feature space to 24,509 unique codes. To ensure a more comparable baseline to the other models, which implicitly incorporate patient context, we explicitly added normalized age at prediction time and coded sex as additional features to the feature matrix used for the count-based baseline.
@@
All approaches were evaluated using a five-fold cross-validation. Within each fold, hyperparameters were optimized on a dedicated validation set of similar sample size to that of the training partition in that fold. For each fold, patients were randomly assigned to train, validation, and test sets. The test set for each task within a fold was constructed by selecting all available positive cases (capped at 10,000), and an equal number of negative cases after allocating patients to the train and validation sets. To specifically assess performance in low-data scenarios, mimicking potential clinical applications with limited examples, we again evaluated models using incrementally larger subsets of the training data, with the final evaluation point capped at a sample size of 256 patients from the training set, even when larger test sets were available.
4.11 Ablation Studies of EHR Serialization
To better understand the contribution of various components in the EHR serialization process to the performance of the LLM embedding models, we conducted a series of ablation studies. These ablations used Qwen3-Emb-8B with a 4,096-token limit due to computational constraints. We first examined the role of task-specific instructions by replacing them with a generic prompt (see Table 9) or removing instructions entirely, thereby isolating their contribution to the resulting embeddings. Next, we systematically excluded individual elements of the serialization, demographics, body metrics, vital signs, laboratory values, visit headings, conditions, medications, and procedures, while keeping all other parts of the pipeline fixed to avoid confounding. Finally, we constructed serializations that retained only one of these components at a time to assess its standalone predictive value. Together, these experiments identify which elements are most critical for producing effective EHR representations and clarify how the models leverage structured clinical data for downstream prediction tasks.
@@ 4.12 Effect of Different Time Windows
To examine the influence of recency on predictive performance, we varied the time window preceding the prediction date used during EHR text serialization. We evaluated Qwen3-Emb-8B, Qwen2-Emb-7B, Llama3.1-LLM2Vec-8B, and the count-based baseline with a GBM head across seven intervals: one hour, one day, one week (7 days), one month (30 days), one year (365 days), three years (1,095 days), and full history. For each window, only events occurring within the specified interval before the prediction time were included. Hence, only data from the respective time window contributed to the aggregated information and the visit data in the EHR serialization. All other aspects of the serialization, including structure, formatting, and instruction prompts remained unchanged to isolate the effect of the temporal window.
@@ 4.13 Effect of Different Context Sizes
We investigated the impact of varying context sizes in the LLM embedding models. We sought to determine whether encoding information from older visits enhances prediction performance and whether longer inputs might dilute critical details, such as laboratory values, in the final embeddings. Specifically, we evaluated Qwen3-Emb-8B, Qwen2-Emb-7B, and Llama3.1-LLM2Vec-8B models with input token limits of 512, 1,024, 2,048, 4,096 and 8,192 tokens. Input tokens exceeding these thresholds were discarded. Due to the design of our EHR serialization process, additional input tokens primarily consisted of medical concepts from past visits. All other preprocessing choices and task-specific instructions were held fixed across context-size settings. By testing these varying context sizes, we aimed to assess the balance between capturing historical medical data and preserving the clarity of high-priority information within the embeddings.
@@ A.1 Additional Experimental Details
| Additional Prompt Added for Decoder Qwen3-8B | |
| Decoder Prompt | Answer STRICTLY with a single token: Yes or No. No punctuation, no extra words. |
A.2 Full Results on EHRSHOT
@@
| Baselines [44] | |||||
| CLMBR-T-Base | |||||
| Count-based + GBM | |||||
| LLM Embedding Models | |||||
| Qwen3-Emb-8B | |||||
| Qwen3-Emb-4B | |||||
| Qwen3-Emb-0.6B | |||||
| Qwen2-Emb-7B | |||||
| Qwen2-Emb-1.5B | |||||
| Llama3.1-LLM2Vec-8B | |||||
| LLM Embedding Model + EHR Foundation Model [44] | |||||
| Qwen3-Emb-8B + CLMBR-T-Base | |||||
| LLM Embedding Model + GBM Head | |||||
| Qwen3-Emb-8B + GBM | |||||
| Multiple Embedding Model for EHR (MEME) [34] with Linear Head | |||||
| Qwen3-Emb-8B MEME | |||||
| MedBERT MEME | |||||
| Encoder Language Models with Chunked Inputs | |||||
| ClinicalBERT | |||||
| MedBERT | |||||
| DeBERTaV3 large | |||||
| DeBERTaV3 base | |||||
| BERT large | |||||
| BERT base | |||||
| EHR Serialization Formats | |||||
| Markdown (ours) | |||||
| List events | |||||
| List events old | |||||
| List events with time | |||||
| List events old with time | |||||
| JSON | |||||
| XML | |||||
| YAML | |||||
| Original Markdown Serialization with 4,096 Tokens | |||||
| Full EHR | |||||
| Instruction Experiments | |||||
| Generic Instruction | |||||
| Empty Instruction | |||||
| Removing Serialization Components | |||||
| No Demographics | |||||
| No Body Metrics | |||||
| No Vital Signs | |||||
| No Lab Results | |||||
| No Visits | |||||
| No Conditions | |||||
| No Medications | |||||
| No Procedures | |||||
| Individual Serialization Components | |||||
| Only Demographics | |||||
| Only Body Metrics | |||||
| Only Vital Signs | |||||
| Only Lab Results | |||||
| Only Visits | |||||
| Only Conditions | |||||
| Only Medications | |||||
| Only Procedures | |||||
| Qwen3-Emb-8B | |||||
| 8,192 context size | |||||
| 4,096 context size | |||||
| 2,048 context size | |||||
| 1,024 context size | |||||
| 512 context size | |||||
| Qwen2-Emb-7B | |||||
| 8,192 context size | |||||
| 4,096 context size | |||||
| 2,048 context size | |||||
| 1,024 context size | |||||
| 512 context size | |||||
| Llama3.1-LLM2Vec-8B | |||||
| 8,192 context size | |||||
| 4,096 context size | |||||
| 2,048 context size | |||||
| 1,024 context size | |||||
| 512 context size | |||||
| Qwen3-Emb-8B (4,096 tokens) | |||||
@@
| 1 hour | |||||
| Llama3.1-LLM2Vec-8B (4,096 tokens) | |||||
| Full patient history | |||||
| 3 years (1,095 days) | |||||
| 1 year (365 days) | |||||
| 1 month (30 days) | |||||
| 1 week (7 days) | |||||
| 1 day | |||||
| 1 hour | |||||
| Count-based + GBM | |||||
| Full patient history | |||||
| 3 years (1,095 days) | |||||
| 1 year (365 days) | |||||
| 1 month (30 days) | |||||
| 1 week (7 days) | |||||
| 1 day | |||||
| 1 hour | |||||
A.3 Additional Performance Results on EHRSHOT
A.4 Additional Performance Results on UK Biobank
| Baselines [44] | ||||
| CLMBR-T-Base | ||||
| Count-based + GBM | ||||
| LLM Embedding Models | ||||
| Qwen3-Emb-8B | ||||
| Qwen2-Emb-7B | ||||
| Llama3.1-LLM2Vec-8B | ||||
A.5 Additional Performance For Encoder and Decoder Models